Making On-Line Presentations with L^{A}TEX

Stuart C. Shapiro Department of Computer Science

State University of New York at Buffalo

226 Bell Hall Buffalo, NY 14260-2000

         shapiro@cs.buffalo.edu

Displaying Slides

Login to local workstation/X-terminal. Execute xhost + Create xterm for file-server — rlogin

Execute setenv DISPLAY ⟨ local ⟩:0.0
Use xdvi or ghostview

Preparing Slides

Use L^{A}TEX

\documentclass{slides}
\begin{document}
      .
      .
      .
\begin{slide}
\begin{center}
<Title>
\end{center}
\vfill
<Contents>
\vfill
\end{slide}
      .
      .
      .
\end{document}

See L^{A}TEX Manual, Section 5.2

Files for Previewing

Source: ⟨file⟩.tex

For xdvi: latex ⟨file⟩.tex ⟶ ⟨file⟩.dvi
For ghostview: texprint -D ⟨file⟩.dvi > ⟨file⟩.ps

Rules for Slides

Short sentences/phrases

Rules for Slides

Short sentences/phrases Notes for speaker & listeners

Rules for Slides

Short sentences/phrases Notes for speaker & listeners Don’t crowd, clutter, over-fill

Rules for Slides

Short sentences/phrases Notes for speaker & listeners Don’t crowd, clutter, over-fill Say everything on slide

Rules for Slides

Short sentences/phrases Notes for speaker & listeners Don’t crowd, clutter, over-fill Say everything on slide Don’t digress

Rules for Slides

Short sentences/phrases Notes for speaker & listeners Don’t crowd, clutter, over-fill Say everything on slide Don’t digress Don’t cover

Rules for Slides

Short sentences/phrases Notes for speaker & listeners Don’t crowd, clutter, over-fill Say everything on slide Don’t digress Don’t cover — use multiple slides instead

Rules for Slides

Short sentences/phrases Notes for speaker & listeners Don’t crowd, clutter, over-fill Say everything on slide Don’t digress Don’t cover — use multiple slides instead Time your slides